# Discontinuous integro-differential equations and sliding mode control

### v2

@@ 3.3 Uniqueness of Filippov solutions of IDEs
$\begin{split}\tfrac{d\|e(t)\|^{2}}{dt}=&2e^{\top}(t)(f_{1}(t)-f_{2}(t))\\ &+\!2\!\!\int^{t}_{t^{0}}\!\!\!e^{\top}\!(t)\Phi(t,\tau)\tilde{f}(\tau,x_{1}(\tau)\!-\!\tilde{f}(\tau,x_{2}(\tau))d\tau\\ \leq&2\ell(t)\|e(t)\|^{2}+2\bar{\Phi}\|e(t)\|\int^{t}_{t^{0}}\tilde{\ell}(\tau)\|e(\tau)\|d\tau\\ \leq&(2\ell(t)+\bar{\Phi}^{2})\|e(t)\|^{2}+\|\tilde{\ell}\|_{L^{2}}^{2}\int^{t}_{t^{0}}\|e(\tau)\|^{2}d\tau,\end{split}\vskip-2.84526pt$ (35)
where $f_{1}(t)\in K[f](t,x_{1}(t)),f_{2}(t)\in K[f](t,x_{2}(t))$. One-Sided Lipschitz Condition for $K[f]$ [Filippov1988:Book, page 106], Lipschitz Condition for $\tilde{f}$ and Cauchy–Schwartz inequality in $L^{2}$ have been utilized on the last steps. Since, $e(t^{0})=0$ then the obtained inequality implies e(t)=0e(t)\!=\!0 for $t\!>\!t^{0}$ (see [Pachpatte1975:JMAA]). ∎
For discontinuous $\tilde{f}$, the uniqueness analysis is challenging.
Let us denote $F_{\mathcal{S}}(t,x,\gamma)\!=\!\left(K[f](t,x)+\gamma\right)\,\cap\,\mathcal{C}_{\mathcal{S}}(x)$, where $\gamma\in\mathbb{R}^{n}$ and $\mathcal{S}\subset\mathbb{R}^{n}$ is the discontinuity set of $f$.
@@
then then the IVP (27), (12) has a Filippov solution being unique, at least, locally in the forward time.
@@
Let us study the case $x^{0}\in\mathcal{S}$. This means that $x^{0}\in\partial G^{j}$ for all $j\in\mathcal{N}(x^{0})$, where, as before, $\mathcal{N}:\mathbb{R}^{n}\to 2$ is the indicator function (i.e., $j\in\mathcal{N}(x_{0})$ if $x_{0}\in\overline{G}^{j}$). Since the function $f$ is piece-wise continuous then there exists $\bar{\epsilon}>0$ such that $\mathcal{N}(x^{0}+\delta)\subset\mathcal{N}(x^{0})$ for $x_{0}+\delta\in S:\|\delta\|\leq\bar{\epsilon}$.
@@
$f^{j}(t^{0}+s,x^{0}+B(\epsilon^{j}))+B(\epsilon^{j})\subset\left(\mathbb{R}^{n}\backslash{\mathcal{C}}_{\bar{G}^{j}}(x^{0})\right)\vskip-2.84526pt$
for all $s\in(0,\epsilon^{j})$. Let $\epsilon=\min\left\{\bar{\epsilon},{\hat{\epsilon}},\min_{j\in\mathcal{N}(x^{0})}\{\epsilon^{j}\}\right\}$. Repeating considerations of the proof of Theorem 8, we conclude that any solution $x^{\epsilon}$ of the IVP
$\dot{x}^{\epsilon}(t)\in K[f](t,x^{\epsilon}(t))+B(\epsilon),\quad x^{\epsilon}(t^{0})=x^{0}\in\mathcal{S}\vskip-2.84526pt$ (36)
belongs to $\mathcal{S}$ as long as $x^{\epsilon}(t)\in x^{0}+B(\epsilon)$. For any continuous function $t\in[t^{0},t^{0}+\epsilon]\to x(t)\in\mathbb{R}^{n}$, the set-valued function
@@ A Filippov solution of the IVP
$\left\{\begin{split}\dot{x}_{1}(t)=&\;u(x(t)),\quad\quad\quad\quad\quad\quad\quad\quad\quad x_{1}(0)=0,\\ \dot{x}_{2}(t)=&\int^{t}_{0}{\sin(t+\tau)}u(x(\tau))d\tau,\quad\quad x_{2}(0)=0,\\ u(x)=&-\mathrm{sign}(x_{1}),\quad\quad\quad\;\;x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}\end{split}\right$. (37)
@@
As a straightforward consequence of the two above theorems we deduce the following corollary.
@@
Remark 1 (On explicit Euler method for IDEs).
$\begin{split}x_{k+1}=&x_{k}+hf(t^{0}+kh,x_{k})+hI_{k},\\ I_{k}=&h\sum^{k}_{i=0}\Phi(t^{0}+kh,t^{0}+ih)\tilde{f}(t^{0}+ih,x_{i}),\vskip-2.84526pt\end{split}$ (38)
(being uniformly bounded and equicontinuous on $\mathcal{I}$) converges to $x$ uniformly on $\mathcal{I}$ as $h\to 0$. This conclusion remains valid if the discontinuous IDE (27) has no sliding mode. In the sliding mode case, similarly to ODEs, the explicit Euler method correctly approximates the Filippov solution of the IDE away an $\epsilon$-neighborhood of a sliding set $\mathcal{S}$, where $\epsilon\to 0$ as $h\to 0$. A consistent discrete-time approximation of the motion on the sliding surface requires development of advanced numerical schemes [AcaryBrogliato2010:SCL], [Polyakov_etal2019:SIAM_JCO]. This goes out of the scope of the paper.
@@ 3.4 Equivalent control method for systems modeled by IDEs
Similarly to ODEs, the set of Utkin solutions of the first kind is a subset of the set of the Utkin solutions of the second kind, since $K[u](t,x(t))\subset(K[u_{1}](t,x(t)),...,K[u_{m}(t,x(t))])^{\top}$. They always coincide for single input systems (i.e., if $m=1$). Utkin solutions always exist for affine-in-control IDEs.
@@
is unique and given by $x_{1}(t)=x_{2}(t)=0,\forall t\geq 0$. The set of Filippov solutions of the IVP (40) coincides with the sets of Utkin solutions of the second kind of the IVP (40) and coincides with the set of Filippov solutions of the IVP (37) (see, Example 5). If control inputs $u_{1}$ and $u_{2}$ are assumed to be independent (e.g., realized by different actuators), then, in practice, the input signals cannot be identical. In this particular case, the Utkin solution of the second kind (40) provides a more realistic description of the system motion than the Utkin solution of the first kind.
@@
    $(\mathcal{K}\tilde{u})(t)\!=\!\!\int^{t}_{t^{*}}\!\!\!G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\tilde{u}(\tau)d\tau,\vskip-2.84526pt$ (42)
@@
    $g(t)\!=\!-G^{-1}(t,x(t))\left.\tfrac{\partial s(y)}{\partial y}\right|_{y=x(t)}\!\left(a(t,x(t))\!+\!\tilde{g}(t)\right),\vskip-5.69054pt$
    $\tilde{g}(t)\!=\!\int^{t^{*}}_{t^{0}}\!\!\!\!\!\Phi(t,\tau)\left(a(\tau,x(\tau))\!+\!b(\tau,x(\tau))u(t,x(\tau)\right)d\tau.\vskip-2.84526pt$
@@
    If $a,b,\tilde{a},\tilde{b},\frac{\partial s}{\partial x}$ are locally Lipschitz continuous and $G(t,y_{1})=G(t,y_{2}),\tilde{b}(t,y_{1})=\tilde{b}(t,y_{2})$, $\forall y_{1},y_{2}\in\mathcal{S}$ then any Utkin solution of (8), (12) belonging to $\mathcal{S}$ is unique.
@@
$\begin{split}\dot{x}(t)\!=&a(t,x(t))+b(t,x(t))u_{\rm eq}(t)\\ &+\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\left(\tilde{a}(\tau,x(\tau))\!+\!\tilde{b}(\tau,x(\tau))u_{\rm eq}(\tau)\right)d\tau,\vskip-5.69054pt\end{split}$ (43)
@@
$u_{\rm eq}(t)\!=\!g(t)\!+\!(\mathcal{K}u_{\rm eq})(t),\quad t\in[t^{*},t^{1}].\vskip-2.84526pt$ (44)
@@
$M\!=\!\!\operatorname*{ess\,sup}_{t^{*}\leq\tau\leq t\leq t^{1}}\|G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\|d\tau\!<\!+\infty.\vskip-5.69054pt$
@@
[OReganMeehan1998:Book] contin. in $(x,u)$ continuous in $x$
@@
in $(x,u)$ smooth sliding surface
@@ 4.1 SMC for a distributed input delay system
where, $t>t_{0}\in\mathbb{R},x(t)\in\mathbb{R}^{n}$, $u(t)\in\mathbb{R}^{m}$, $\Phi:\Delta\to\mathbb{R}^{n\times n}$ be locally essentially bounded on $\Delta$, $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R}^{m})$ is a matched perturbation, $p\in L^{\infty}(\mathbb{R};\mathbb{R}^{n})$ is a possibly mismatched perturbation and $A\in\mathbb{R}^{n\times n},B,\tilde{B}\in\mathbb{R}^{n\times m}$ are known matrices. This IDE can be considered as a linear system with a distributed input delay. SMC design for a linear system with a distributed state delay can be found in [Gouaisbaut_etal2001:TDS]. The aim is to stabilize the output
@@
to zero in a finite time. For ODEs (i.e. for $\Phi=\mathbf{0}$), the SMC theory (see, [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book]) suggests the feedback
$u(t)=-\rho(t)(CB)^{-1}\frac{y(t)}{\|y(t)\|_{P}},\quad\rho(t)>0,\vskip-2.84526pt$ (47)
where $\|y\|_{P}=\sqrt{y^{\top}Py}$ is the weighted Euclidean norm with a positive definite matrix $P\succ 0$. Let $\|Q\|_{P}=\sup_{y\neq 0}\frac{\|Qy\|_{P}}{\|y\|_{P}}$, $Q\in\mathbb{R}^{m\times m}$ be the matrix norm is induced by the weighted Euclidean norm $\|\cdot\|_{P}$. The condition $\det(CB)\neq 0$ means that the output has the relative degree 1 with respect to the input. This is the standard assumption for the first order sliding mode control design [Utkin1992:Book], [Utkin_etal2020:Book]. If the output has a higher relative degree with respect to the input, then, to solve the considered regulation problem, the high order sliding mode control design algorithms [Levant2003:IJC], [Shtessel_etal2014:Book] must be extended to the IDE (45). This extension goes out of the scope of the paper.
@@
[…]
$M=\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|C\Phi(t,\tau)\tilde{B}(CB)^{-1}\|_{P}d\tau<1,\vskip-2.84526pt$ (49)
$\rho(t)\!\geq\!\tfrac{\left\|C(B\gamma(t)+p(t))\right\|_{P}+\int^{t}_{t^{0}}\!\|C\Phi(t,\tau)\tilde{B}\gamma(\tau)\|_{P}d\tau+\delta}{1-M},\;\delta\!>\!0,\vskip-2.84526pt$ (50)
@@
$\begin{split}\tfrac{d\|x(t)\|}{dt}\!\leq&\|Ax(t)\|\!+\!\rho\|B(CB)^{-1}y\|/\|y\|_{P}\!+\!\|B\gamma\|_{L^{\infty}}\!+\!\|p\|_{L^{\infty}}\!\\ &\!\!+\!\left(\!\rho\|\tilde{B}(CB)^{-1}y\|/\|y\|_{P}\!+\!\|\tilde{B}\gamma\|_{L^{\infty}}\!\right)\!\!\int^{t}_{t^{0}}\!\!\|\Phi(t,\tau)\|d\tau.\end{split}\vskip-2.84526pt$
Since $\Phi$ is measurable and locally bounded on $\Delta$ then the function $t\mapsto\int^{t}_{t^{0}}\|\Phi(t,\tau)\|d\tau$ is measurable and locally bounded on $t_{0},+\infty)$. So, using Grönwall–Bellman inequality, we conclude that $\|x(t)\|<+\infty$ for any finite $t\geq t^{0}$.
@@
$\begin{split}\dot{y}(t)\in&\,{\Lambda y(t)+}CB\gamma(t)-\tfrac{\rho y(t)}{\|y(t)\|_{P}}+Cp(t)\\ &+\!\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}\left(\gamma(\tau)-\tfrac{\rho(CB)^{-1}y(\tau)}{\|y(\tau\|)_{P}}\right)d\tau\end{split}\vskip-2.84526pt$ (52)
@@
The above theorem presents a sufficient condition under which $Cx=0$ is a sliding surface of the closed-loop system (45), (46), (47). The SMC theory [Utkin1992:Book] also studies a motion of the system on this surface.
Similarly to ODEs, the sliding motion can be studied using equivalent control method.
@@
$p=\mathbf{0}$ and $\Phi(t,\tau)=\xi(t-\tau)I_{n}$, where $\xi(t)=1$ for $0\leq t-\tau<1$ and $\xi(t)=0$ for $t-\tau\geq 1$. In this case, the condition (49) is fulfilled with $M=1$ and $|CB|=2$. Therefore, by Theorem 20, the SMC (47) with $\rho>6\|\gamma\|_{L^{\infty}}$ stabilizes the output $y=Cx$ of the system (45) to zero in a finite time. The results of numerical simulations for $\rho=4$, $\gamma(t)=0.5\cos(2t)$ and $t^{0}=0$ are depicted on Figures 1, 2. The simulations have been done by means of explicit Euler method (38) with the sampling period $h=0.001$. The integral is computed using the method of rectangles using the same sampling period.
The control signal depicted on Figure 3 is the typical signal generated by SMC, which has a bounded chattering during the sliding mode (after the time instant $t=T\approx 0.55$). Recall that (see, e.g., [Utkin1992:Book]), in practice, the equivalent control can be estimated by filtering the discontinuous input signal $u(t)$:
@@
can be considered as a sliding mode indicator. It must be close to zero for $t\geq T$, provided that Utkin solution is an adequate description of a motion of the considered discontinuous control system. The evolution of the signal $\delta$ computed for $\epsilon=0.01$ is depicted on Figure 4. The numerical computation of $\delta$ has been done using the sampling period $h=0.0001$. The obtained $\delta$ is, indeed, close to zero for $t\geq 0.6$. This indirectly confirms a validity of the Utkin solutions for description of a motion of discontinuous SMC control systems modeled by IDEs. A small delay in detection of SMC is caused by the inertia of the linear filter (62) and by numerical errors.

### v3

@@ 3.3 Uniqueness of Filippov solutions of IDEs
$\small\begin{split}\!\!\!\!\!\tfrac{d\|e(t)\|^{2}}{dt}\!=&2e^{\!\top}\!(t)\!\left(\!f_{1}(t)\!-\!f_{2}(t)\!+\!\!\!\int^{t}_{t^{0}}\!\!\Phi(t,\tau)\left(\tilde{f}_{1}(\tau)\!-\!\tilde{f}_{2}(\tau)\right)d\tau\!\right)\\[-5.69054pt] \leq&2\ell(t)\|e(t)\|^{2}+2\bar{\Phi}\|e(t)\|\int^{t}_{t^{0}}\tilde{\ell}(\tau)\|e(\tau)\|d\tau\\[-5.69054pt] \leq&2\ell(t)\|e(t)\|^{2}\!+\!(\bar{\Phi}\|e(t)\|)^{2}\!+\!\left(\int^{t}_{t^{0}}\!\!\tilde{\ell}(\tau)\|e(\tau)\|d\tau\right)^{2}\\[-5.69054pt] \leq&2\ell(t)\|e(t)\|^{2}+\bar{\Phi}^{2}\|e(t)\|^{2}+\|\tilde{\ell}\|_{L^{2}}^{2}\!\!\int^{t}_{t^{0}}\!\!\|e(\tau)\|^{2}d\tau,\end{split}\!\!\!\!\vskip-2.84526pt$ (36)
where $f_{1}(t)\in K[f](t,x_{1}(t)),f_{2}(t)\in K[f](t,x_{2}(t))$, $\tilde{f}_{1}(\tau)=\tilde{f}(\tau,x_{1}(\tau))$, $\tilde{f}_{2}(\tau)=\tilde{f}(\tau,x_{2}(\tau))$, $\overline{\Phi}\geq 0$ is defined in the proof of Theorem 14. One-Sided Lipschitz Condition for $f$ (and, consequently, for $K[f]$, see [17, page 106]) with $\ell\in L^{\infty}_{\rm loc}$, Lipschitz Condition for $\tilde{f}$ with $\tilde{\ell}\in L^{\infty}_{\rm loc}$ and Cauchy–Schwartz inequality in $L^{2}((0,t);\mathbb{R})$ have been utilized to derive (36). Since, $e(t^{0})=0$ then the obtained inequality implies e(t)=0e(t)\!=\!0 for $t\!>\!t^{0}$ (see [39]). ∎
Let $F_{\mathcal{S}}(t,x,\gamma)\!=\!\left(K[f](t,x)+\gamma\right)\,\cap\,\mathcal{C}_{\mathcal{S}}(x)$, where $\gamma\in\mathbb{R}^{n}$, $\mathcal{S}\!\subset\!\mathbb{R}^{n}$ is discontinuity set of $f$ and $\mathcal{C}_{\mathcal{S}}$ is tangent cone for $\mathcal{S}$.
@@
Let us study the case $x^{0}\in\mathcal{S}$. This means that $x^{0}\in\partial G^{j}$ for all $j\in\mathcal{N}(x^{0})$, where, as before, $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ is the indicator function, i.e., $j\in\mathcal{N}(x_{0})$ if $x_{0}\in\partial G^{j}$. Since the function $f$ is piece-wise continuous then there exists $\bar{\epsilon}>0$ such that $\mathcal{N}(x^{0}+\delta)\subset\mathcal{N}(x^{0})$ for $x_{0}+\delta\in S:\|\delta\|\leq\bar{\epsilon}$.
@@
$f^{j}(t^{0}+s,x^{0}+\mathcal{B}(\epsilon^{j}))+\mathcal{B}(\epsilon^{j})\subset\left(\mathbb{R}^{n}\backslash{\mathcal{C}}_{\bar{G}^{j}}(x^{0})\right),\;\forall s\in(0,\epsilon^{j}),\vskip-2.84526pt$
where $\mathcal{B}(\epsilon^{j})\subset\mathbb{R}^{n}$ is an open ball of the radius $\epsilon^{j}>0$.
Let $\epsilon=\min\left\{\bar{\epsilon},\hat{\epsilon},\min_{j\in\mathcal{N}(x^{0})}\{\epsilon^{j}\}\right\}$. Repeating the proof of Theorem 8, we conclude that any solution $x^{\epsilon}$ of the IVP
$\dot{x}^{\epsilon}(t)\in K[f](t,x^{\epsilon}(t))+\mathcal{B}(\epsilon),\quad x^{\epsilon}(t^{0})=x^{0}\in\mathcal{S}\vskip-2.84526pt$ (37)
belongs to $\mathcal{S}$ as long as $x^{\epsilon}(t)\in x^{0}+\mathcal{B}(\epsilon)$. For any continuous function $t\in[t^{0},t^{0}+\epsilon]\to x(t)\in\mathbb{R}^{n}$, the set-valued function
@@ A Filippov solution of the IVP
$\vskip-2.84526pt\left\{\begin{split}\dot{x}_{1}(t)=&\;u(x(t)),\quad\quad\quad\quad\quad\quad\quad\quad\quad x_{1}(0)=0,\\[-2.84526pt] \dot{x}_{2}(t)=&\int^{t}_{0}\sin(t+\tau)u(x(\tau))d\tau,\quad\quad x_{2}(0)=0,\\[-2.84526pt] u(x)=&-\mathrm{sign}(x_{1}),\quad\quad\quad\;\;x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}\end{split}\right$. (38)
@@
The corollary given below follows from above theorems.
@@
Theorems 17, 18 and Corollary 2 prove only a local-in-time uniqueness of solutions. For ODEs, due to the semigroup property of solutions (see, e.g., [10]), the local-in-time uniqueness implies the uniqueness on the whole interval of the existence of solution. For IDEs, the semigroup property is not valid, so, a Filippov solution of IDE may be unique on a short interval of time and non-unique on a long interval of time. Utkin solutions introduced below solve this problem in some cases.
@@ 3.4 Equivalent control method for systems modeled by IDEs
Similarly to ODEs, the set of Utkin solutions of the first kind is a subset of the set of the Utkin solutions of the second kind, since $K[u](t,x(t))\subset(K[u_{1}](t,x(t)),...,K[u_{m}](t,x(t)))^{\top}$. They always coincide for single input systems (i.e., if $m=1$). Utkin solutions always exist for affine-in-control IDEs.
@@
is unique and given by $x_{1}(t)=x_{2}(t)=0,\forall t\geq 0$. The set of Filippov solutions of the IVP (40) coincides with the sets of Utkin solutions of the second kind of the IVP (40) and coincides with the set of Filippov solutions of the IVP (38) (see, Example 5). If control inputs $u_{1}$ and $u_{2}$ are assumed to be independent (e.g., realized by different actuators), then, in practice, the input signals cannot be identical. In this particular case, the Utkin solution of the second kind provides a more realistic description of the system motion than the Utkin solution of the first kind.
Utkin solutions of the first kind for an affine-in-control IDE with a smooth sliding surface may be unique globally in time.
@@
    $\small(\mathcal{K}\tilde{u})(t)\!=\!\!\int^{t}_{t^{*}}\!\!\!G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\tilde{u}(\tau)d\tau,\vskip-2.84526pt$ (42)
@@
    $\small\!\!g(t)\!=\!-G^{-1}(t,x(t))\left.\tfrac{\partial s(y)}{\partial y}\right|_{y=x(t)}\!\left(a(t,x(t))\!+\!\tilde{g}(t)\right),\vskip-5.69054pt$
    $\small\tilde{g}(t)\!=\!\int^{t^{*}}_{t^{0}}\!\!\!\!\!\Phi(t,\tau)\left(a(\tau,x(\tau))\!+\!b(\tau,x(\tau))u(t,x(\tau)\right)d\tau.\vskip-2.84526pt$
@@
    If $a,b,\tilde{a},\tilde{b},\frac{\partial s}{\partial x}$ are locally Lipschitz continuous and
    $G(t,y_{1})=G(t,y_{2}),\tilde{b}(t,y_{1})=\tilde{b}(t,y_{2}),\quad\forall y_{1},y_{2}\in\mathcal{S}$
    then any Utkin solution of (8), (12) belonging to $\mathcal{S}$ is unique.
@@
$\small\begin{split}\dot{x}(t)\!=&a(t,x(t))+b(t,x(t))u_{\rm eq}(t)\\[-2.84526pt] &+\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\left(\tilde{a}(\tau,x(\tau))\!+\!\tilde{b}(\tau,x(\tau))u_{\rm eq}(\tau)\right)d\tau,\vskip-5.69054pt\end{split}$ (43)
@@
$\small u_{\rm eq}(t)\!=\!g(t)\!+\!(\mathcal{K}u_{\rm eq})(t),\quad t\in[t^{*},t^{1}].\vskip-2.84526pt$ (44)
@@
$\small M\!=\!\!\operatorname*{ess\,sup}_{t^{*}\leq\tau\leq t\leq t^{1}}\|G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\|d\tau\!<\!+\infty.\vskip-5.69054pt$
@@
[34] contin. in $(x,u)$ continuous in $x$
@@
in $(x,u)$ smooth and invariant
sliding surface
Remark 1 (On explicit Euler method for IDEs).
$\vskip-2.84526pt\begin{split}x_{k+1}=&x_{k}+hf(t^{0}+kh,x_{k})+hI_{k},\\ I_{k}=&h\sum^{k}_{i=0}\Phi(t^{0}+kh,t^{0}+ih)\tilde{f}(t^{0}+ih,x_{i}),\end{split}$ (45)
(being uniformly bounded and equicontinuous on $\mathcal{I}$) converges to $x$ uniformly on $\mathcal{I}$ as $h\to 0$. This conclusion remains valid if the discontinuous IDE (28) has no sliding mode.
In the sliding mode case, similarly to ODEs, the explicit Euler method correctly approximates the Filippov/Utkin solution of the IDE away an $\epsilon$-neighborhood of a sliding set $\mathcal{S}$, where $\epsilon\to 0$ as $h\to 0$. A consistent discrete-time approximation of the motion on the sliding surface requires development of advanced numerical schemes [1], [42]. This goes out of the scope of the paper.
@@ 4.1 SMC for a distributed input delay system
where, $t>t_{0}\in\mathbb{R},x(t)\in\mathbb{R}^{n}$, $u(t)\in\mathbb{R}^{m}$, $\Phi:\Delta\to\mathbb{R}^{n\times n}$ be locally essentially bounded on $\Delta$, $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R}^{m})$ is a matched perturbation, $p\in L^{\infty}(\mathbb{R};\mathbb{R}^{n})$ is a possibly mismatched perturbation and $A\in\mathbb{R}^{n\times n},B,\tilde{B}\in\mathbb{R}^{n\times m}$ are known matrices. This IDE can be considered as a linear system with a distributed input delay. SMC design for a linear system with a distributed state delay can be found in [20]. The aim is to stabilize the output
@@
to zero in a finite time. For ODEs (i.e. for $\Phi=\mathbf{0}$), the SMC theory (see, [49], [14], [46]) suggests the feedback
$u(t)=-\rho(CB)^{-1}\frac{y(t)}{\|y(t)\|_{P}},\quad\rho>0,\vskip-2.84526pt$ (48)
where $\|y\|_{P}=\sqrt{y^{\top}Py}$ is the weighted Euclidean norm with a positive definite matrix $P\succ 0$. Let $\|Q\|_{P}=\sup_{y\neq 0}\frac{\|Qy\|_{P}}{\|y\|_{P}}$, $Q\in\mathbb{R}^{m\times m}$ be the matrix norm is induced by the weighted Euclidean norm $\|\cdot\|_{P}$. The condition $\det(CB)\neq 0$ means that the output has the relative degree 1 with respect to the input. This is the standard assumption for the first order sliding mode control design [49], [50]. If the output has a higher relative degree with respect to the input, then, to solve the considered regulation problem, the high order sliding mode control design algorithms [29], [46] must be extended to the IDE (46). This extension goes out of the scope of the paper.
@@
$\Lambda^{\top}P+P\Lambda\preceq 0,\quad 0\prec P=P^{\top}\in\mathbb{R}^{m\times m},\vskip-2.84526pt$ (49)
$\rho\!\geq\!\tfrac{\left\|C(B\gamma(t)+p(t))\right\|_{P}+\int^{t}_{t^{0}}\!\|C\Phi(t,\tau)\tilde{B}\gamma(\tau)\|_{P}d\tau+\delta}{1-M},\vskip-2.84526pt$ (50)
$\small\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|C\Phi(t,\tau)\tilde{B}(CB)^{-1}\|_{P}d\tau\leq M<1,\;\;\delta\!>\!0,\vskip-2.84526pt$ (51)
@@
$\small\begin{split}\tfrac{d\|x(t)\|}{dt}\!\leq&\|Ax(t)\|\!+\!\rho\|B(CB)^{-1}y\|/\|y\|_{P}\!+\!\|B\gamma\|_{L^{\infty}}\!+\!\|p\|_{L^{\infty}}\!\\ &\!\!+\!\left(\!\rho\|\tilde{B}(CB)^{-1}y\|/\|y\|_{P}\!+\!\|\tilde{B}\gamma\|_{L^{\infty}}\!\right)\!\!\int^{t}_{t^{0}}\!\!\|\Phi(t,\tau)\|d\tau.\end{split}\vskip-2.84526pt$
Since $\Phi$ is measurable and locally bounded on $\Delta$ then the function
$t\mapsto\int^{t}_{t^{0}}\|\Phi(t,\tau)\|d\tau$
is measurable and locally bounded on $t_{0},+\infty)$. So, using Grönwall–Bellman inequality, we conclude that $\|x(t)\|<+\infty$ for any finite $t\geq t^{0}$.
@@
$\small\begin{split}\dot{y}(t)\in&\,\Lambda y(t)+CB\gamma(t)-\tfrac{\rho y(t)}{\|y(t)\|_{P}}+Cp(t)\\ &+\!\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}\left(\gamma(\tau)-\tfrac{\rho(CB)^{-1}y(\tau)}{\|y(\tau\|)_{P}}\right)d\tau\end{split}\vskip-2.84526pt$ (53)
@@
If $\|\gamma\|_{L^{\infty}}\leq\bar{\gamma}$ and $\|p\|_{L^{\infty}}\leq\bar{p}$ then the inequality (51) holds for $\rho=(\sqrt{\|P\|}(\|CB\|\bar{\gamma}+\|C\|\bar{p}+M\|CB\|\bar{\gamma})+\delta)/(1-M)$. Such a design based on upper estimates on uncertain parameters of the system perfectly follows SMC theory [49], [14] and appreciated by SMC practice [48].
The above theorem presents a sufficient condition under which $Cx=0$ is a sliding surface of the closed-loop system (46), (47), (48). The SMC theory [49] also studies a motion of the system on this surface. Similarly to ODEs, the sliding motion can be studied using equivalent control method.
@@
Hence, the condition (61) implies that $\operatorname*{ess\,sup}_{s\geq t}\|\tilde{u}_{\rm eq}(t)\|\to 0$ as $t\to+\infty$ and
$\operatorname*{ess\,sup}_{s\geq t}\left\|\int^{t}_{t-d}\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau\right\|\to 0\text{ as }t\to+\infty.$
@@
$p=\mathbf{0}$ and $\Phi(t,\tau)=\xi(t-\tau)I_{n}$, where $\xi(t)=1$ for $0\leq t-\tau<1$ and $\xi(t)=0$ for $t-\tau\geq 1$. In this case, the condition (51) is fulfilled with $M=1$ and $|CB|=2$. Therefore, by Theorem 20, the SMC (48) with $\rho>6\|\gamma\|_{L^{\infty}}$ stabilizes the output $y=Cx$ of the system (46) to zero in a finite time. The results of numerical simulations for $\rho=4$, $\gamma(t)=0.5\cos(2t)$ and $t^{0}=0$ are depicted on Figures 1, 2 and 3. The simulations have been done by means of explicit Euler method (45) with the sampling period $h=0.001$. The integral is computed using the method of rectangles using the same sampling period. The applicability of this method is discussed in Remark 1.
To compare the SMC for IDE with the SMC for ODE , we also present the simulation results of the system (46)-(48) with $\Phi=\mathbf{0}$ (please see Figures 4, 5 and 6). The simulations show a faster convergence of the ODE ($\Phi=\mathbf{0}$) system to zero.
The control signal depicted on Figure 3 is the typical signal generated by SMC, which has a chattering during the sliding mode (after the time instant $t=T\approx 0.55$). The chattering is caused by the use of the explicit method, which correctly approximates the dynamics of the SMC system only away from the sliding surface (see Remark 1 or [1]). Recall that (see, e.g., [49]), in practice, the equivalent control can be estimated by filtering the discontinuous input signal $u(t)$:
@@
can be considered as a practical sliding mode indicator. Indeed, if the Utkin solution is an adequate description of a motion of the considered discontinuous control system and the explicit Euler method gives a precise enough approximate of Utkin solution of IDE then the approximate identity $\sigma(t)\approx 0$ may be treated as an indication of a practical sliding mode for $t\geq T$. The evolution of the signal $\delta$ computed for $\epsilon=0.01$ is depicted on Figure 7. The numerical computation of $\delta$ has been done with the sampling period $h=0.0001$. The obtained $\delta(t)$ is, indeed, close to zero for $t\geq 0.6$. This may be interpreted as some heuristic confirmation of the validity of Utkin solutions for description of a motion of discontinuous SMC control systems modeled by IDEs. A small delay in detection of SMC is caused by an inertia of the linear filter (63) and by numerical errors of the explicit Euler method. Development of a rigorous numerical analysis for IDEs with SMC is the important problem for the future research.